OpenStudy (anonymous):

can somebody post the different formulas i can use to find the slope of lines?

OpenStudy (anonymous):

Like point slope, slope intercept?

OpenStudy (anonymous):

yea like those. thers like 4 of them i think

OpenStudy (anonymous):

and the one for two points and stuff like that

OpenStudy (anonymous):

Point slope, y-y1=m(x-x1) Slope intercept, y=mx+b those are the two main ones. What others do you need?

OpenStudy (anonymous):

\[m = \frac{y2 -y1}{x2 - x1}\]

OpenStudy (anonymous):

Ah yeah, thats how you find slope.

OpenStudy (anonymous):

I think there are only three. -Point-Slope -Slope Intercept -And definition of slope